3.Given that x + 2y + 3z =1, 3x + 2y + z = 4, x + 3y + 2z = 0. What is x,y and z?
a.7/4,−3/4,1/4
b.5/4,−2/4,¼
c.\(1,-3/4,1/4)
d.None of the option

4.What is the angle between the two linesâˆ′3x+4y=8 and âˆ′2xâˆ′8yâˆ′14=0?
a.14.63 degree
b.âˆ′14degree
c.−14.63degree
d.14 degree
1
Expert's answer
2019-04-15T11:46:21-0400

Given that x + 2y + 3z =1, 3x + 2y + z = 4, x + 3y + 2z = 0, Cramer's rule:


"x=\frac{\begin{Vmatrix}
 1 & 2 & 3 \\
 4 & 2 & 1 \\
0 & 3 & 2
\end{Vmatrix}}
{\begin{Vmatrix}
 1 & 2 & 3 \\
 3 & 2 & 1 \\
1 & 3 & 2
\end{Vmatrix}} = \frac{21}{12}=\frac{7}{4}\,\,""y=\frac{\begin{Vmatrix}
 1 & 1 & 3 \\
 3 & 4 & 1 \\
 1 & 0 & 2
\end{Vmatrix}}
{\begin{Vmatrix}
 1 & 2 & 3 \\
 3 & 2 & 1 \\
1 & 3 & 2
\end{Vmatrix}} = \frac{-9}{12}=\frac{-3}{4}\,\,"

"z=\frac{\begin{Vmatrix}
 1 & 2 & 1 \\
 3 & 2 & 4 \\
 1 & 3 & 0
\end{Vmatrix}}
{\begin{Vmatrix}
 1 & 2 & 3 \\
 3 & 2 & 1 \\
1 & 3 & 2
\end{Vmatrix}} = \frac{3}{12}=\frac{1}{4}"

Solution: a. "7/4, \, -3/4, \, 1/4"

What is the angle between the two lines "a\cdot x+b \cdot y=c" and "d\cdot x + e\cdot y =f"

Vectors, perpendicular to lines: "\vec{v}_1 = (a,b), \, \vec{v}_2 = (d,e)"

Angle between vectors:

"\cos{\phi}= \frac{\vec{v}_1\cdot \vec{v}_2}{|\vec{v}_1|\cdot|\vec{v}_2|} = \frac{ad+eb}{\sqrt{a^2+b^2}\sqrt{d^2+e^2}}"

If "\cos{\phi}<0 \Rightarrow \phi> \pi/2" , and angle between lines equal to "\pi - \phi" .

Therefore, angle between lines:


"\phi=\arccos \left| \frac{ad+eb}{\sqrt{a^2+b^2}\sqrt{d^2+e^2}}\right|"